# Identifying diffusive length scales in one-dimensional Bose gases

### 12 Jan 2024 v2

@@ 2 Propagation of quasi-particles in the 1D Bose gas
By virtue of integrability, the 1D Bose gas can be solved exactly using the Bethe Ansatz, which identifies the elementary excitations as fermionic quasi-particles uniquely labelled by their quasi-momentum, or rapidity, $\theta$ [32, 33]. In the thermodynamics limit, the density of occupied rapidities, i.e. the density of quasi-particles, $\rho_{\mathrm{p}}(\theta)$ fully characterises the thermodynamic properties of the local equilibrium macrostate. Similarly, a density of holes $\rho_{\mathrm{h}}(\theta)$ can be introduced, which describes the density of unoccupied rapidities; together the two densities form the density of states $\rho_{\mathrm{s}}(\theta)$. The occupational fraction of allowed rapidity states, dubbed the filling function $\vartheta(\theta)=\rho_{\mathrm{p}}(\theta)/\rho_{\mathrm{s}}(\theta)$, equivalently characterises the macrostate.
At thermal equilibrium, the state of a homogeneous 1D Bose gas with density $n$ is characterized by only two dimensionless parameters: the interaction strength $\gamma=c/n$ and the reduced temperature $\mathcal{T}=\frac{2\hbar^{2}k_{\mathrm{B}}T}{mg^{2}}$, where $T$ is the temperature and $k_{\mathrm{B}}$ Boltzmann’s constant. The interaction strength and temperature span an entire phase-diagram of the 1D Bose gas, whose regimes can be identified via the $g^{(2)}$-function [35]. At asymptotic values of $\gamma$ and $\mathcal{T}$, the description of the Bose gas drastically simplifies. However, in order to describe the state at arbitrary values of these parameters, the Thermodynamic Bethe Ansatz (TBA) is generally needed [36]. For more detail on the TBA and GHD of the 1D Bose gas, see Appendix A and Refs. [37, 38, 39]).
@@ 2.1 Generalized Hydrodynamics at the diffuse scale
where $\Delta(\theta,\theta^{\prime})=-\frac{2c}{c^{2}+(\theta-\theta^{\prime})^{2}}$ is the two-body scattering kernel of the Lieb-Liniger model. Further, in eq. (2) we have employed bold symbols to indicate integral operators and the following shorthand notation for their action
@@
2.2 Linearization of the diffusive GHD equation
@@
The single-mode picture above has clear analogies to that of phonons in Bogoliubov theory for quasi-condensates [48] (or more generally in Luttinger liquid theory [49, 50]): For these effective low-energy Hamiltonians, collective phase-density excitations in the form of phonons are the exact eigenstates. In analog quantum field simulators, basis of phonons have been used extensively to analyze experimental results (see e.g. Refs. [27, 28, 29, 30, 31]). However, when considering the underlying microscopic Hamiltonian (here the Lieb-Liniger Hamiltonian), phonons are superpositions of the true eigenstates and therefore do not have infinite lifetime. In Ref. [51], the phonon lifetime was derived by expressing the phonon as a coherent superposition of Bethe particle-hole states; over time, dephasing of the particle-hole states leads to an apparent relaxation of the phonon, similarly to the filling of a single $k$-mode $\delta\vartheta_{j}$.
2.3 Critical length scale of diffusion
@@
3 Magnitude of diffusive corrections in thermal states
Figure 2: Propagation metric $\Gamma$ scaled by the coupling strength $c$ for thermal states at different dimensionless temperatures $\mathcal{T}$ and interaction strengths $\gamma$. Lower values signify a relative increased contribution of diffusion to quasi-particle propagation. The solid lines indicate the boundary between the three main thermodynamic regimes of the Lieb-Liniger model, namely the quasi-condensate (QC), ideal Bose gas (IBG), and and Tonks-Girardeau gas (TG). The dashed lines mark additional sub-regimes, namely the thermal-quantum statistics boundary for the fluctuations of the quasi-condensate and the degeneracy transition of the ideal Bose gas.
For the purpose of quantifying the contribution of diffusion to the overall propagation of quasi-particle in the state $\rho_{\mathrm{p}}$, we introduce the measure
$\Gamma=\frac{1}{n}\int_{-\infty}^{\infty}\mathrm{d}\theta\>\rho_{\mathrm{p}}(\theta)\frac{|v^{\mathrm{eff}}(\theta)|}{D^{\mathrm{diag}}(\theta)}\;,$ (19)
where $D^{\mathrm{diag}}(\theta)=w(\theta)\rho_{\mathrm{s}}^{-2}(\theta)$ are the diagonal elements of the diffusion kernel and $n$ the atomic density. The ratio ${|v^{\mathrm{eff}}(\theta)|}/{D^{\mathrm{diag}}(\theta)}$ describes the distance travelled by a particle with rapidity $\theta$ relative to the variance of its trajectory [13]. The measure $\Gamma$ is a weighted mean of said ratio with respect to the density of quasi-particles, thus estimating the overall relation between ballistic and diffusive propagation in the state $\rho_{\mathrm{p}}$. For fixed values of $\gamma=c/n$ we find that $\Gamma$ scales linearly with the coupling strength $c$. Hence, we calculate $\Gamma/c$ for a large number of thermal states for various combinations of reduced temperatures $\mathcal{T}$ and interaction strength $\gamma$. The results are plotted in figure 2, where the different regimes of the Lieb-Liniger phase diagram have been indicated by solid lines.
((a))
((b))
((c))
Finally, it should be stressed that $\Gamma$ can be computed for any state $\vartheta$, not just thermal states. Indeed, explicitly constructing a state which maximizes fluctuations will result in a larger diffusive broadening of the quasi-particle trajectory. An important non-thermal state, is the initial state of the seminal quantum Newton’s cradle experiment [54]; here, two halves of a (typically thermal) state have been boosted to large, opposite rapidities [55]. Notably, boosting the state substantially increases the quasi-particle velocities while merely shifting $D^{\mathrm{diag}}(\theta)$ towards higher rapidities, thus resulting in an increase in $\Gamma$. Nevertheless, diffusive effects are important to the long time-scale dynamics of the quantum Newton’s cradle, as purely ballistic propagation can produce quasi-particle distributions featuring very fine structures in the ($z,\theta$)-phase-space [56]. Such small wavelength features are much more susceptible to diffusive effects, which over time lead to the features being washes out and the system thermalizing [46].
Diffusive effects become increasingly important at shorter length scales. To quantify these length scales, we can calculate the critical mode $j^{\star}$ of eq. (18). Note that $\Gamma$ and $j^{\star}$ are closely related; states with low $\Gamma$ will likewise feature lower critical modes. One particular system, where this sort of analysis is very insightful, is the single-mode quench in a quasi-condensate realized in the recent experiment of Ref. [24]. Although the Bose gas was realized in a parameter regime where diffusive effects are rather weak (based on our previous analysis), the ability to excite a single $k$-mode (of the atomic density) makes such an experimental platform ideal for studying diffusive effects.
4.1 Experimental setup and protocol
In the experiment, a Bose gas of ${}^{87}$b atoms are trapped on an atom chip [57], which creates a strong magnetic potential along two (transverse) axes, while providing weak trapping along the third (longitudinal or 1D) axis, effectively realizing a quasi-1D gas. A digital micromirror device (DMD) enables the creation of arbitrary optical potentials along the 1D axis [58]; for this protocol, the atoms are initially trapped in a box potential of length $L=80\>\mu\mathrm{m}$. By modulating the bottom of the box trap in the shape of a cosine, a density perturbation in the shape of single $k$-mode is imprinted in the condensate (see Fig. 4(a)). Once confined in the box, the condensate density is 60-80 $\text{atoms}/\mu\mathrm{m}$, and the initial state is well-described by a thermal state with temperature in the range 50-120 nK (depending on the degree of cooling performed). Here we consider three quenches with dimensionless interaction strength and temperature:
  1. (a)
    $\gamma=1.8\cdot 10^{-3}$ and $\mathcal{T}=1.2\cdot 10^{3}$,
  2. (b)
    $\gamma=2.1\cdot 10^{-3}$ and $\mathcal{T}=2.9\cdot 10^{3}$,
  3. (c)
    $\gamma=1.6\cdot 10^{-3}$ and $\mathcal{T}=2.8\cdot 10^{3}$.
Comparing with figure 2, we find that the gas is realized in the quasi-condensate phase near the thermal-quantum statistics boundary for the fluctuations; in this parameter regime, diffusive effects are expected to be weak.
To initiate dynamics, the confining potential is quenched to a flat-bottomed box trap at time $t=0$ and the subsequent dynamics and relaxation is monitored by recording the atomic density via absorption imaging. An example of the measured density evolution is plotted in figure 4(b). Owing to a high degree of control over the 1D potential, a single density mode of the condensate can be excited, that is, the atomic density inside of the box trap is accurately given by $n(z,t)=n_{0}+\delta n_{j}(t)\cos\left(k_{j}z\right)$, where $\delta n_{j}(t)$ is the amplitude of the mode and $k_{j}=2\pi j/L$ is its momentum. For the quenches analysed here, only the $j=1$ mode is excited, however, the ability to also address higher modes was also demonstrated in Ref. [24]. In the context of Bogoliubov theory, the excited density mode corresponds to an eigenstate of the effective low-energy Hamiltonian [48]. However, as previously mentioned, the phononic basis does not represent the true eigenstates of the system [51]. Further, due to the high temperature of the condensate, the applicability of the low-energy model is very limited; we will discuss and demonstrate this in the following section.
((a))
((b))
((c))
4.2 Analysis of quasi-particle propagation
4.2.1 Comparing experimental observations with linearized GHD
((a))
((b))
((c))
((d))
((e))
To assess whether a linearization of the dynamics is valid for the system, we simulate the dynamics of the gas using linearized GHD [47] and compare with experimental observations. First, to extract the filling function of the background and perturbation, we spatially Fourier transform the filling of the initial thermal states and identify the $j=0$ mode as the stationary background. See figure 4(c) for an example. The perturbation $\delta\vartheta(z,\theta)$ is comprised of all remaining modes; since density is a non-linear function of the filling, an excitation of a single-density mode generally features a perturbation $\delta\vartheta$ containing multiple modes. Indeed, after Fourier transforming $\delta\vartheta(z,\theta)$ we find higher modes of the filling to be occupied, however, their amplitude is relative low (less than 10% of the $j=1$ mode), whereby we will focus solely on the evolution of $\delta\vartheta_{j=1}(t,\theta)$.
4.2.2 Quasi-particle dephasing and applicability of phonon basis
With contributions of both diffusion and a dimensional crossover being negligible, the apparent relaxation of the density mode seen in figure 5 must be due to the ballistic dephasing of quasi-particle trajectories. In figure 5(e) we plot the coefficients $\eta_{\omega,j=1}$, representing the population of each propagation eigenstate for the $j=1$ perturbation. The eigenstates are ordered by the imaginary part (ballistic contribution) of their corresponding eigenvalue $\alpha_{\omega,j=1}$. At low values of $\alpha_{\omega,j=1}$ the population of the corresponding eigenstates is zero, as these states belongs to the background $\vartheta_{0}(\theta)$. For all quenches, the population $\eta_{\omega,j=1}$ is peaked around the edge of the background and features long thermal tails. We find that $\eta_{\omega,j}$ is mostly independent of the mode number $j$; instead its width/extend depends on the temperature, while its total area depends on the initial mode amplitude. Thus, one can immediately understand the temperature dependence of the relaxation rate: A larger spread in populated ballistic eigenvalues $\alpha_{\omega,j=1}$ leads to a faster dephasing of the eigenstates comprising the mode $\delta\vartheta_{j=1}$. In the context of phonons, the variation of the populations $\eta_{\omega,j=1}$ is thus an indication of the phonon lifetime. Given the fast relaxation of the density mode (particularly in the hot realizations), a phonon basis is not particularly suitable for describing the system. However, as the temperature decreases, the population of large $\alpha_{\omega,j=1}$ vanishes, substantially increasing the phonon lifetime. For this particular setup, a condensate of 30 nK (half the temperature of quench (a)) would be reasonably described by an effective low-energy Hamiltonian.
4.2.3 Diffusive length scales
@@
Finally, we calculate the critical mode $j_{\omega}^{\star}$, whose spectra are shown in figure 5(d), also labelled by the imaginary part of the corresponding eigenvalue. We find that $j_{\omega}^{\star}\gg 1$ for all eigenstates, demonstrating that the GHD dynamics of the addressed mode is entirely dominated by ballistic propagation of the quasi-particles. From the spectrum of $j_{\omega}^{\star}$ we may identify the length scales at which diffusion would become relevant. First, we note that the critical mode is highly dependent on the eigenstate: Since $\alpha_{\omega,j=1}$ denotes the ballistic contribution to quasi-particle propagation, we would expect $j_{\omega}^{\star}$ to be minimal around low $\alpha$-values. This is indeed the case, however, the corresponding states all belong to the stationary background. Likewise, at large $\alpha$ the critical mode is very high. Interestingly, at intermediate values of $\alpha_{\omega,j=1}$, the spectra of $j_{\omega}^{\star}$ all feature a local minimum. The location of the minimum is related to the states with maximal thermal thermal fluctuations (see eq. (9)). Further, comparing with the eigenstate population of the perturbation of figure 5(e), we find that the location of the minimum coincides with the most populated eigenstates; the observed dynamics of the system is thus mainly driven by the evolution of these particular eigenstates. From figure 5(d), we find a minimum critical mode of $j_{\omega}^{\star}\sim 200$. Hence, given the same experimental parameters, diffusive effects would become dominant for quenches of the $j=200$ mode and higher. For an experimental box length of $L=80$ µm, the corresponding diffusive length scale is around 0.4 µm, which is close to the healing length of the condensate. At such length scales, higher order (e.g. dispersive) corrections become relevant for dynamics [60, 61, 62, 63]. Further, from an experimental perspective, length scales on that order are typically not resolvable by the available imaging techniques.

### 12 Mar 2024 v3

@@ 2 Propagation of quasi-particles in the 1D Bose gas
By virtue of integrability, the 1D Bose gas can be solved exactly using the Bethe Ansatz, which identifies the elementary excitations as fermionic quasi-particles uniquely labelled by their quasi-momentum, or rapidity, $\theta$ [32, 33]. In the thermodynamics limit, the density of occupied rapidities, i.e. the density of quasi-particles, $\rho_{\mathrm{p}}(\theta)$ fully characterises the thermodynamic properties of the local equilibrium macrostate. Similarly, a density of holes $\rho_{\mathrm{h}}(\theta)$ can be introduced, which describes the density of unoccupied rapidities; together the two densities form the density of states $\rho_{\mathrm{s}}(\theta)$. The occupational fraction of allowed rapidity states, dubbed the filling function $\vartheta(\theta)=\rho_{\mathrm{p}}(\theta)/\rho_{\mathrm{s}}(\theta)$, equivalently characterises the macrostate. The quasi-particle density of a thermal state at a given coupling strength and temperature is obtained by solving the Thermodynamic Bethe Ansatz (TBA) [35] (see also Appendix A and Refs. [36, 37] for more details).
Finally, the large scale dynamics of the system can be described by means of Generalized Hydrodynamics (GHD) [8, 9], which formulates a hydrodynamic theory of integrable models in terms of the propagation of its Bethe Ansatz quasi-particles. In the following section 2.1, we briefly review the main concepts of quasi-particle propagation and the diffusive GHD equations of the 1D Bose gas. For more details, see the reviews [38, 39]. Next, in section 2.2, we quantify the diffusive contribution to quasi-particle propagation in thermal states.
@@ 2.1 Generalized Hydrodynamics at the diffuse scale
where $\Delta(\theta,\theta^{\prime})=\frac{-2c}{c^{2}+(\theta-\theta^{\prime})^{2}}$ is the two-body scattering kernel of the Lieb-Liniger model. Further, in eq. (2) we have employed bold symbols to indicate integral operators and the following shorthand notation for their action
@@
2.2 Quantifying operator spreading in thermal states
According to Generalized Hydrodynamics, operator spreading in integrable systems is driven by the propagation of the Bethe Ansatz quasi-particles. For conserved charges in particular, each quasi-particle carries a certain amount of charge density specified by the corresponding single-particle eigenvalue of the charge, which generally is a function of rapidity. After traveling through a homogeneous system for a duration $t$, a quasi-particle with rapidity $\theta$ will ballistically have propagated a distance $\bar{z}(t,\theta)=t|v^{\mathrm{eff}}(\theta)|$, while the diffusive broadening of the quasi-particle trajectory is [13]
$\delta\bar{z}^{2}(t,\theta)$ $=t\rho_{\mathrm{s}}^{-2}(\theta)\int_{\infty}^{\infty}\mathrm{d}\theta^{\prime}\rho_{\mathrm{p}}(\theta^{\prime})\left(1-\vartheta(\theta^{\prime})\right)\left[\frac{\Delta^{\mathrm{dr}}(\theta,\theta^{\prime})}{2\pi}\right]^{2}\left|v^{\mathrm{eff}}(\theta)-v^{\mathrm{eff}}(\theta^{\prime})\right|$ (10)
$=t\rho_{\mathrm{s}}^{-2}(\theta)w(\theta)\;.$
For the fastest quasi-particle of the system, $\bar{z}$ and $\delta\bar{z}^{2}$ coincide with the position and diffusive broadening of the operator spreading front, respectively. Thus, in order to quantify the contribution of diffusion to the overall propagation of quasi-particles in the state $\rho_{\mathrm{p}}$, we introduce the measure
$\Gamma$ $=\frac{1}{n}\int_{-\infty}^{\infty}\mathrm{d}\theta\>\rho_{\mathrm{p}}(\theta)\frac{\bar{z}(\theta)}{\delta\bar{z}^{2}(\theta)}$ (11)
$=\frac{1}{n}\int_{-\infty}^{\infty}\mathrm{d}\theta\>\rho_{\mathrm{p}}(\theta)\rho_{\mathrm{s}}^{2}(\theta)\frac{|v^{\mathrm{eff}}(\theta)|}{w(\theta)}\;,$
where $n=\int_{-\infty}^{\infty}\mathrm{d}\theta\>\rho_{\mathrm{p}}(\theta)$ is the atomic density. Physically, $\Gamma$ describes the average distance travelled by quasi-particles relative to the variance of their trajectory, weighted according to the distribution of particles.
At thermal equilibrium, the state of a homogeneous 1D Bose gas is characterized by just two dimensionless parameters: the interaction strength $\gamma=c/n$ and the reduced temperature $\mathcal{T}=\frac{2mk_{\mathrm{B}}T}{\hbar^{2}c^{2}}$, where $T$ is the temperature and $k_{\mathrm{B}}$ Boltzmann’s constant. These two parameters span an entire phase-diagram of the 1D Bose gas, whose regimes can be identified via the $g^{(2)}$-function [45]. For thermal states with fixed values of $\gamma$, we find that $\Gamma$ scales linearly with the coupling strength $c$. Therefore, $\gamma$ and $\mathcal{T}$ fully determine the scaled measure $\Gamma/c$ for thermal states, which we compute for a large span of parameters and plot in figure 2. For comparison, the different regimes of the Lieb-Liniger phase diagram have been indicated by solid lines.
Figure 2: Propagation metric $\Gamma$ scaled by the coupling strength $c$ for thermal states at different dimensionless temperatures $\mathcal{T}$ and interaction strengths $\gamma$. Lower values signify a relative increased contribution of diffusion to quasi-particle propagation. The solid lines indicate the boundary between the three main thermodynamic regimes of the Lieb-Liniger model, namely the quasi-condensate (QC), ideal Bose gas (IBG), and Tonks-Girardeau gas (TG). The dashed lines mark additional sub-regimes, namely the thermal-quantum statistics boundary for the fluctuations of the quasi-condensate and the degeneracy transition of the ideal Bose gas.
((a))
((b))
((c))
Finally, it should be stressed that $\Gamma$ can be computed for any state $\vartheta$, not just thermal states. Indeed, explicitly constructing a state which maximizes fluctuations will result in a larger diffusive broadening of the quasi-particle trajectory. An important non-thermal state, is the initial state of the seminal quantum Newton’s cradle experiment [48]; here, two halves of a (typically thermal) state have been boosted to large, opposite rapidities [49]. Notably, boosting the state substantially increases the quasi-particle velocities while merely shifting $w(\theta)$ towards higher rapidities, thus resulting in an increase in $\Gamma$. Nevertheless, diffusive effects are important to the long time-scale dynamics of the quantum Newton’s cradle, as purely ballistic propagation can produce quasi-particle distributions featuring very fine structures in the ($z,\theta$)-phase-space [50]. Such small wavelength features are much more susceptible to diffusive effects, which over time lead to the features being washes out and the system thermalizing [51].
The measure $\Gamma$ describes the relative contribution of ballistic and diffusive quasi-particle propagation in a given state. However, the measure fails to account for the role of inhomogeneity; indeed, diffusive effects become increasingly important at shorter length scales. In this section, we employ a linearized version of the diffusive GHD equation to analyze the dynamics of a single momentum mode of the atomic density, i.e. an excitation with a well-defined length scale. Such a single-mode quench was achieved in the recent experiment of Ref. [24]. Although the Bose gas was realized in the quasi-condensate regime where diffusive effects are very weak, the ability to excite a single momentum mode makes such an experimental platform ideal for studying dynamics at different length scales.
3.1 Experimental setup and protocol
In the experiment, a Bose gas of ${}^{87}$b atoms are trapped on an atom chip [52], which creates a strong magnetic potential along two (transverse) axes, while providing weak trapping along the third (longitudinal or 1D) axis, effectively realizing a quasi-1D gas. A digital micromirror device (DMD) enables the creation of arbitrary optical potentials along the 1D axis [53]; for this protocol, the atoms are initially trapped in a box potential of length $L=80\>\mu\mathrm{m}$. By modulating the bottom of the box trap in the shape of a cosine, a density perturbation in the shape of single momentum mode is imprinted in the condensate (see Fig. 4(a)). Once confined in the box, the condensate density is 60-80 $\text{atoms}/\mu\mathrm{m}$, and the initial state is well-described by a thermal state with temperature in the range 50-120 nK (depending on the degree of cooling performed). Here we consider three quenches with dimensionless interaction strength and temperature:
  1. (a)
    $\gamma=1.8\cdot 10^{-3}$ and $\mathcal{T}=1.2\cdot 10^{3}$,
  2. (b)
    $\gamma=2.1\cdot 10^{-3}$ and $\mathcal{T}=2.9\cdot 10^{3}$,
  3. (c)
    $\gamma=1.6\cdot 10^{-3}$ and $\mathcal{T}=2.8\cdot 10^{3}$.
Both the coupling strength and temperature of the system are acquired by independent measurement; the former is computed from the transverse trapping frequency, while the latter is measured by means of density ripples thermometry [54, 55, 56]. Comparing with figure 2, we find that the gas is realized in the quasi-condensate phase near the thermal-quantum statistics boundary for the fluctuations; in this parameter regime, diffusive effects are expected to be weak.
To initiate dynamics, the confining potential is quenched to a flat-bottomed box trap at time $t=0$ and the subsequent dynamics and relaxation is monitored by recording the atomic density via absorption imaging. An example of the measured density evolution is plotted in figure 4(b). Owing to a high degree of control over the 1D potential, a single density mode of the condensate can be excited, that is, the atomic density inside of the box trap is accurately given by $n(z,t)=n_{0}+\delta n_{j}(t)\cos\left(k_{j}z\right)$, where $\delta n_{j}(t)$ is the amplitude of the mode and $k_{j}=2\pi j/L$ is its momentum. For the quenches analysed here, only the $j=1$ mode is excited, however, the ability to also address higher modes was demonstrated in Ref. [24]. In the context of Bogoliubov theory, the excited density mode corresponds to an eigenstate of the effective low-energy Hamiltonian [57]. However, as previously mentioned, the phononic basis does not represent the true eigenstates of the system [58]. Further, due to the high temperature of the condensate, the applicability of the low-energy model is very limited; we will discuss and demonstrate this in the following section.
((a))
((b))
((c))
3.2 Analysis of the single-mode dynamics
3.2.1 Linearization of the diffusive GHD equation
@@
3.2.2 Comparing experimental observations with linearized GHD
((a))
((b))
((c))
((d))
((e))
To assess whether a linearization of the dynamics is valid for the experimental system, we simulate the dynamics of the gas using the linearized GHD [60] and compare with experimental observations. First, to extract the filling function of the background and perturbation, we spatially Fourier transform the filling of the initial thermal states and identify the $j=0$ mode as the stationary background. See figure 4(c) for an example. The perturbation $\delta\vartheta(z,\theta)$ is comprised of all remaining modes; since density is a non-linear function of the filling, an excitation of a single-density mode generally features a perturbation $\delta\vartheta$ containing multiple modes. Indeed, after Fourier transforming $\delta\vartheta(z,\theta)$ we find higher modes of the filling to be occupied, however, their amplitude is relative low (less than 10% of the $j=1$ mode), whereby we will focus solely on the evolution of $\delta\vartheta_{j=1}(t,\theta)$.
3.2.3 Quasi-particle dephasing and applicability of phonon basis
With contributions of both diffusion and a dimensional crossover being negligible, the apparent relaxation of the density mode seen in figure 5 must be due to the ballistic dephasing of quasi-particle trajectories. In figure 5(e) we plot the coefficients $\eta_{\omega,j=1}$, representing the population of each propagation eigenstate for the $j=1$ perturbation. The eigenstates are ordered by the imaginary part (ballistic contribution) of their corresponding eigenvalue $\alpha_{\omega,j=1}$. At low values of $\alpha_{\omega,j=1}$ the population of the corresponding eigenstates is zero, as these states belongs to the background $\vartheta_{0}(\theta)$. For all quenches, the population $\eta_{\omega,j=1}$ is peaked around the edge of the background and features long thermal tails. We find that $\eta_{\omega,j}$ is mostly independent of the mode number $j$; instead its width/extend depends on the temperature, while its total area depends on the initial mode amplitude. Thus, one can immediately understand the temperature dependence of the relaxation rate: A larger spread in populated ballistic eigenvalues $\alpha_{\omega,j=1}$ leads to a faster dephasing of the eigenstates comprising the mode $\delta\vartheta_{j=1}$.
This picture has clear analogies to that of phonons in Bogoliubov theory for quasi condensates [57] (or more generally in Luttinger liquid theory [62, 63]): For these effective low-energy Hamiltonians, collective phase-density excitations in the form of phonons are the exact eigenstates. In analog quantum field simulators, basis of phonons have been used extensively to analyze experimental results (see e.g. Refs. [27, 28, 29, 30, 31]). However, when considering the underlying microscopic Hamiltonian (here the Lieb-Liniger Hamiltonian), phonons are superpositions of the true eigenstates and therefore do not have infinite lifetime. In Ref. [58], the phonon lifetime was derived by expressing the phonon as a coherent superposition of Bethe particle-hole states; over time, dephasing of the particle-hole states leads to an apparent relaxation of the phonon, similarly to the filling of a single $k$-mode $\delta\vartheta_{j}$.
In the context of phonons, the variation of the populations $\eta_{\omega,j=1}$ is thus an indication of the phonon lifetime. Given the fast relaxation of the density mode (particularly in the hot realizations), a phonon basis is not particularly suitable for describing the system. However, as the temperature decreases, the population of large $\alpha_{\omega,j=1}$ vanishes, substantially increasing the phonon lifetime. For this particular setup, a condensate of 30 nK (half the temperature of quench (a)) would be reasonably described by an effective low-energy Hamiltonian, even at long timescales.
3.2.4 Calculating diffusive length scales
@@
Calculating the critical mode $j_{\omega}^{\star}$ for the experimentally realized quasi-condensates, we obtain the spectra shown in figure 5(d). We find that $j_{\omega}^{\star}\gg 1$ for all eigenstates, again demonstrating that the GHD dynamics of the addressed mode is entirely dominated by ballistic propagation of the quasi-particles. From the spectrum of $j_{\omega}^{\star}$ we can identify the length scales at which diffusion would become relevant: First, we note that the critical mode is highly dependent on the eigenstate. Since $\alpha_{\omega,j=1}$ denotes the ballistic contribution to quasi-particle propagation, we would expect $j_{\omega}^{\star}$ to be minimal around low $\alpha$-values. This is indeed the case, however, the corresponding states all belong to the stationary background. Likewise, at large $\alpha$ the critical mode is very high. Interestingly, at intermediate values of $\alpha_{\omega,j=1}$, the spectra of $j_{\omega}^{\star}$ all feature a local minimum. The location of the minimum is related to the states with maximal thermal thermal fluctuations (see eq. (9)). Further, comparing with the eigenstate population of the perturbation of figure 5(e), we find that the location of the minimum coincides with the most populated eigenstates; the observed dynamics of the system is thus mainly driven by the evolution of these particular eigenstates. From figure 5(d), we find a minimum critical mode of $j_{\omega}^{\star}\sim 200$. Hence, given the same experimental parameters, diffusive effects would become dominant for quenches of the $j=200$ mode and higher. For an experimental box length of $L=80$ µm, the corresponding diffusive length scale is around 0.4 µm, which is close to the healing length of the condensate. At such length scales, higher order (e.g. dispersive) corrections become relevant for dynamics [64, 65, 66, 67]. Further, from an experimental perspective, length scales on that order are typically not resolvable by the available imaging techniques.